An investment pays simple interest, and quadruples in 13 years. What is the interest rate?
Answer by Theo(13342)

the formula for simple interest is i * p * n.

i is the interest rate per time period.
p is the principal.
n is the number of time periods.

the formula for future value of an investment using simple interest is f = p + p * i * n).

this can be simplified to f = p * (1 + i * n).

you are given that n = 13 years.
you will be looking for the interest rate per year.
since you will be quadrupling your money, than f will be equal to 4p.

your formula of f = p * (1 + i * n) becomes:

4p = p * (1 + 13 * i)

divide both sides of this equation by p to get:

4 = 1 + 13 * i

subtract 1 from both sides of this equation to get:

3 = 13 * i

divide both sides of this equation by 13 to get:

3/13 = i

that's your interest rate per year.

i = 3/13 = .2307692308

for example:

assume your principal is equal to 500 dollars.

your interest in 13 years will be equal to 500 * 3/13 * 13 which will become 500 * 3 which will become 1500 dollars.

add that to your principal of 500 dollars and you get a total of 2000 dollars which is equal to 4 times the original investment of 500 dollars.